OpenStudy (anonymous):

Solve each equation and check each solution (x-3)/5+ (X-2)/2=1/2

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge \frac {x-3}{5} + \frac {x-2}{2}=\frac1{2}$$

OpenStudy (anonymous):

yes that`s the question I asked sir

OpenStudy (skullpatrol):

Recall that when adding or subtracting fractions you rewrite the fractions using their least common denominator.

OpenStudy (anonymous):

yes LCD IS 10 Now I am on this step 2x-15+5x-10=5 than what?

OpenStudy (skullpatrol):

$$\Huge 2x-6+5x-10=5$$

OpenStudy (skullpatrol):

not -15

OpenStudy (anonymous):

yea that`s where i made mistake let me see if i can handle from there thanks for your time

OpenStudy (skullpatrol):

Thanks for asking and trying to understand.

OpenStudy (anonymous):

no problem no i get stack

OpenStudy (anonymous):

like terms

OpenStudy (anonymous):

i meant now

OpenStudy (skullpatrol):

yep

OpenStudy (anonymous):

i have 7x-6-10=5 so i ADD 6 on both sides isn`t it

OpenStudy (skullpatrol):

ok

OpenStudy (anonymous):

so i get 7x-4=11 Add 4 it give me 7x=15

OpenStudy (anonymous):

is that how i was supposed to do

OpenStudy (skullpatrol):

wait...your -10 turned into a -4 for no reason

OpenStudy (anonymous):

ok please show me how you would finish

OpenStudy (anonymous):

2x-6+5x-10=5

OpenStudy (skullpatrol):

$$\Huge 7x-6-10=5$$ $$\Huge 7x-16=5$$

OpenStudy (anonymous):

how do you get 16 since you have 2x-6+5x-10=5 you are right but please tell me how you get -16

OpenStudy (skullpatrol):

$$\Huge 7x-6-10=5$$ $$\Huge 7x+(-6)+(-10)=5$$ $$\Huge 7x+(-16)=5$$ $$\Huge 7x-16 =5$$

OpenStudy (skullpatrol):

$$\Huge a-b= a+(-b)$$ $$\Huge \text{To subtract b, add the opposite of b}$$

OpenStudy (anonymous):

ok thank so much you really helpful

OpenStudy (anonymous):

i do appreciate your time

OpenStudy (skullpatrol):

So what is the solution set?

OpenStudy (anonymous):

3 which is what you get sir

OpenStudy (anonymous):

7x-16=5 Add 16 on both sides 7x=21 divide 7x x=3

OpenStudy (skullpatrol):

Yes, now what number makes the equation x = 3 a true statement?

OpenStudy (skullpatrol):

What number would you replace x with to make a true statement?

OpenStudy (anonymous):

7(3)-16 =5 is that ok

OpenStudy (skullpatrol):

Sure, also x = 3 is obviously true when x is replaced by 3, this gives 3 = 3, right?

OpenStudy (anonymous):

yes thank you so much

OpenStudy (skullpatrol):

What you did is ok too, since 21-16 =5 and 5 = 5.

OpenStudy (anonymous):

today was my first day here I really find it helpful

OpenStudy (skullpatrol):

Take the value of 3 and substitute it in the ORIGINAL equation for x, to see if you get a true statement. This is the best way to check for errors :)

OpenStudy (skullpatrol):

Solve each equation and check each solution

OpenStudy (anonymous):

you mean the original where ever i see X put 3

OpenStudy (skullpatrol):

Yes,you are exactly correct.

OpenStudy (anonymous):

ok let me see if I can do it

OpenStudy (anonymous):

Exactly i get it 5=5

OpenStudy (skullpatrol):

Therefore the solution set is {3}.

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

Thank you so much I would let you go I really appreciate spending you precious time

OpenStudy (anonymous):

so where are you from and where do live now?

OpenStudy (anonymous):

ok sir have a good rest of the day it is late at night here I get to hit the hay